OpenStudy (anonymous):

Can someone tell if my table for this problem is right? If it isn't, can you help me fix it? Mr. Forte is using a brand new bus for his route today. When he starts, the bus has only 75 miles on it. He travels for a total of 4 hours at an average speed of 50 miles per hour, making one stop each hour. Create a graph, equation, table, and sentence to represent the total mileage on the bus as a function of hours driven at 50 miles per hour. Let the independent variable (x) be time and the dependent variable (y) be miles the bus has traveled. I know the equation is y=50x + 75.

Directrix (directrix):

>>>Can someone tell if my table for this problem is right? Please post your table. Thanks.

Directrix (directrix):

Does your table look like this: X Y 0 75 1 125 2 175 3 225 4 275

OpenStudy (anonymous):

This is the table that I came up with- __________________ ]substitute value of [x(hours)] x into the equation [ y(miles) ] [ 0 ] y = 50(4) + 75 [ 75 ] [ 1 ] y = 50(4) + 75 [ 125 ] [ 2 ] y = 50(4) + 75 [ 175 ] [ 3 ] y = 50(4) + 75 [ 225 ] [ 4 ] y = 50(4) + 75 [ 275 ]

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

Sorry it took me a bit to post the table because my laptop is jacked up

Directrix (directrix):

In the table, the x values for time are all set for 4. They should progress 0, 1, 2, 3, and 4.

OpenStudy (anonymous):

\(\rlap{\color{red}{\huge\bigstar}\huge \color{green}{ \text{Welcome to Open Study! }}\color{red}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{Welcome to Open Study! }}\color{yellow}\bigstar}\)

OpenStudy (anonymous):

@i_love_bands26

Directrix (directrix):

y=50(0)+75: y=75 y=50(1)+75: y=125 y=50(2)+75: y=175 y=50(3)+75: y=225 y=50(4)+75: y=275

OpenStudy (anonymous):

Ok thank you so much this really helps @Directrix !!!!

Directrix (directrix):

There is more to do so don't stop now.

Directrix (directrix):

We have the graph and sentence to go. What do you mean by your computer is "jacked up?"

OpenStudy (anonymous):

It just doesn't really work with typing stuff on this site.

OpenStudy (anonymous):

what is that graph for?

Directrix (directrix):

@i_love_bands26 There is no computer built that works well with this site. I changed to FoxFire but still have to refresh the page every 3 minutes.

Directrix (directrix):

Create a graph, equation, table, and sentence to represent the total mileage on the bus as a function of hours driven at 50 miles per hour.

OpenStudy (anonymous):

This site sucks other than the math help from people.

OpenStudy (anonymous):

I just have the graph and the sentence left.

OpenStudy (anonymous):

oh that's great thanks @MARC_

OpenStudy (anonymous):

I'm using chrome @Directrix.

Directrix (directrix):

If the driver drove the bus at an average of 50 mph, the total mileage of the bus is equal to the beginning mileage (75 miles) plus 50 miles times the number of hours of travel for the total trip.

Directrix (directrix):

The people here say Chrome is the best browser for this site. I don't know about that.

OpenStudy (anonymous):

Which part @MARC_ ?

Directrix (directrix):

Graph

Directrix (directrix):

That is the graph of the line but you want only the part in the first quadrant because the bus is not going back in time.

OpenStudy (anonymous):

For my problem @Directrix?

OpenStudy (anonymous):

ok

Directrix (directrix):

Well, the posted graph does show the distance traveled by the bus so I think it is okay.

OpenStudy (anonymous):

Thanks @Directrix!!

Directrix (directrix):

You are welcome.